OpenStudy (anonymous):

can anyone help me out Solve for x: 5 − (x + 5) > −2(x + 4)

OpenStudy (anonymous):

distribute first

OpenStudy (michele_laino):

hint: left side is: 5-x-5=...? right side is: -2x-2*4=...?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

distributing??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok ex: 2(4x+4) 1. multiply the 2 times 4x 2. then multiply the 2 times 4 3. basically whatever is beside the number thats beside the parenthesis on the left side is what you are going to multiply on the inside. 4. the ex. was... 8x+8....

OpenStudy (anonymous):

ok but it say to subtract on the equation

OpenStudy (jdoe0001):

how would you solve say \(\large \bf 5 − (x + 5) = −2(x + 4)\) ?

OpenStudy (anonymous):

me??

OpenStudy (anonymous):

or @Dubba2000

OpenStudy (jdoe0001):

hmm nope, @Dubba2000

OpenStudy (anonymous):

idk

OpenStudy (jdoe0001):

?

OpenStudy (jdoe0001):

so.... if you haven't done linear simplification how are you expected to solve it then? seems to me you'd need to cover that first

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i got -8 as the answer would hat be right

OpenStudy (jdoe0001):

so... how did you get -8 though?

OpenStudy (anonymous):

nm i did it wrong

OpenStudy (anonymous):

i added 4 to x and got 4x and then multiplied it by -2